The volume of blood ejected by the humming birds heart per beat is approximately 1cm3. The average heart rate of the humming bird is 1250 beats per minute. Determine the volume of blood that passes through the heart of a humming bird in one year, report your answer in m3/year.
Answer #1

Number of heart beats in 1 minute = 1250

Minutes in one year = 365×24×60 min. = 525600 min.

Number of heart beats in 1 minute = 525600×1250 = 657×106

Volume of blood passes in one beat = 1 cm​​​​​3​​​​ = 1×10-6 m​​​​​​3

( 1m = 100 cm)

Volume of blood passes in one year = 657×106 ×10-6 m​​​​​​3​​​​​/year

= 657 m​​​​​​3​​​​​/year

Please give feedback!